OpenStudy (anonymous):

simplify using a table(-5y+3)(2y+5)

OpenStudy (anonymous):

@praxer

OpenStudy (praxer):

here, (-5y+3)(2y+5)=0 so, either -5y+3 =0 => y= 3/5 or 2y+5=0 => y=-5/2 so y = -5/2 or 3/5

OpenStudy (anonymous):

the answer choices is -10y^2+31y+15 -10y^2-31y-15 -10y^2+19y-15 -10y^2-19y+15

OpenStudy (praxer):

ok then -10y^2-25y+6y+15 =-10y^2-19y+15 option D. is correct.... :)

OpenStudy (anonymous):

show me the work

OpenStudy (praxer):

first multiply -5y both with 2y and +5 respectively and then multiply 3 with 2y and +5 respectively you will get -10y^2-25y+6y+15 now add the y coefficients than you will get. -10y^2-19y+15.

OpenStudy (anonymous):

thank u

OpenStudy (praxer):

welcome...... :)